It takes a train the same time to travel 450 km as it takes a bus to travel 300 km. If the bus is traveling 30 km/h slower than the train, find the rate at which each is traveling.


Pls. ans. this....

Answer by mananth(16946)

train rate = x
bus rate = x-30
..
train distance = 450 miles
bus distance = 300 miles
..
time bus = time train
450/x= 300/(x-30)
cross mulyiply
450(x-30)=300x
450x-13500=300x
450x-300x=13500
150x=13500
/150
x=13500/150
rate of train = 90 mph.
rate of bus = 90-30 = 60 mph
...
m.ananth@hotmail.ca